# Zero-shot Meta-learning for Tabular Prediction Tasks with Adversarially Pre-trained Transformer

### v1

@@ 4.2 Zero-shot Learning
Recent work such as Xian et al. (2018, 2017); Chang et al. (2008); Larochelle et al. (2008); Palatucci et al. (2009) have shown impressive capability of zero-shot learning in the space of language and vision problems. Recent approaches to zero-shot or few-shot learning for tabular data problems mostly encode tabular data as language, and then leverage large language models (LLMs) for their zero- or few-shot capabilities (see Hegselmann et al. (2023); Nam et al. (2023); Gardner et al. (2024)). These approaches rely on relevant information about the tabular data existing in LLMs – this is most obviously the case when column names are meaningful – but it is not guaranteed for all tabular data problems.
@@ 4.3 Adversarial Training
Upon generative adversarial networks (GANs) (Goodfellow et al., 2015; Madry et al., 2018; Kurakin et al., 2017), recent work such as Shafahi et al. (2019) improved on the efficiency by combining the back-propagation steps of the generator and discriminator. However, this method has been shown to suffer from catastrophic overfitting (Andriushchenko & Flammarion, 2020; Kim et al., 2021) without further modifications. Other works focusing on improving the efficiency of GAN training include Wong et al. (2020) and Zhang et al. (2019) where they restrict most of the forward and back propagation within the first layer of the network during adversarial updates. Zhang et al. (2021) in particular noted that weights updates frequently going back and forth in opposite directions in one training epoch, suggest those updates are redundant. Many other variations have been introduced to mitigate vanishing gradient and additional challenges of GAN training (see Jabbar et al. (2021) and references therein): failing at finding a Nash-equilibrium (Ratliff et al., 2016), and internal covariate shift (Ioffe, 2015).
@@ 5 Conclusion
In this work, we gave the first effort in using adversarial synthetic data generators for the pre-training of deep zero-shot meta-learning algorithms. We proposed APT, a zero-shot meta-learner that improves the performance of TabPFN on tabular prediction tasks and matches state-of-the-art GBDTs on small tabular classification tasks. In addition, we proposed a mixture block neural design within the transformer architecture to eliminate the class size restriction of PFNs, addressing a crucial problem on zero-shot classifications. As for limitations, APT does not outperform GBDTs on large tabular datasets, and shared the quadratic runtime and memory scaling of TabPFN. Hence, extensions of this work could explore means of acquiring data representations in a more inexpensive manner. Besides, future research in this field could shift the attention slightly more to zero-shot regressions, where there is a lot room for improvement. The final prediction layer of most neural architecture has reduced dimensions for regression compared to classification, while continuous variables generally have higher empirical cardinality than class labels – such dissonance could drive more creative solutions that benefit both zero-shot and traditional DL pipelines in solving tabular regression problems.
@@ Acknowledgements
We thank Tyler Farnan, Gang Mei and C. Bayan Bruss for the insightful discussions.
@@ Appendix A Background
In this section, we give a brief introduction to PFNs, and specifically the synthetic data generating mechanism of TabPFN. For a more complete description, see Müller et al. (2021); Hollmann et al. (2022); Nagler (2023). Given training dataset $D^{\mathrm{train}}=(X^{\mathrm{train}},\bm{y}^{\mathrm{train}})$ and testing dataset $D^{\mathrm{test}}=(X^{\mathrm{test}},\bm{y}^{\mathrm{test}})$, the goal is to approximate the posterior predictive distribution (PPD) $\bm{y}^{\mathrm{test}}\sim p(\cdot|X^{\mathrm{test}},D^{\mathrm{train}})$. In the Bayesian framework for supervised learning, the prior of the dataset is a hypothesis of the data generating mechanism $\phi$ drawn from hypothesis space $\Phi$, and the PPD can be factorized as
$p(\cdot|X^{\mathrm{test}},D^{\mathrm{train}})$ $=\int_{\phi\in\Phi}p(\cdot|X^{\mathrm{test}},\phi)p(\phi|X^{\mathrm{test}},D^{\mathrm{train}})\,d\phi$ (4)
$=\int_{\phi\in\Phi}p(\cdot|X^{\mathrm{test}},\phi)\frac{p(\phi)p(D^{\mathrm{train}}|\phi)p(X^{\mathrm{test}}|\phi)}{p(X^{\mathrm{test}},D^{\mathrm{train}})}\,d\phi$ (5)
$\propto\int_{\phi\in\Phi}p(\cdot|X^{\mathrm{test}},\phi)p(D^{\mathrm{train}}|\phi)p(X^{\mathrm{test}}|\phi)p(\phi)\,d\phi,$ (6)
PFNs conduct synthetic prior fitting by defining a family of data generating mechanisms $\Phi$ to draw synthetic datasets $D=D^{\mathrm{train}}\cup D^{\mathrm{test}}\sim p(D)=\mathbb{E}_{p(\phi)}[p(D|\phi)]$, and use a transformer model $q_{\theta}(\cdot|X^{\mathrm{test}},D^{\mathrm{train}})$ to approximate $p(\cdot|X^{\mathrm{test}},D^{\mathrm{train}})$. The loss is given by
$\mathcal{L}(q_{\theta})=\mathbb{E}_{p(D)}\left[-\log q_{\theta}(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},D^{\mathrm{train}})\right]$ (7)
@@ A.2 Limitations
Although there is no theoretical limitation on the number of data PFNs can handle, the transformer architecture does entail significant computation complexity and memory usage for large datasets. Besides, given the nature of MLP input embedding layer and MLP final prediction layer, there is a theoretical limitation on the number of features and the number of classes that PFNs can handle. The former is less of an issue since feature selections or simply random sampling of features can be performed, and PFNs would still yield ideal performance as shown in McElfresh et al. 2024. The latter is a rather big problem for classification tasks because there is hardly any direct and effective work-around.
Appendix B Hyperparameter Settings
@@
Appendix C More Results

### v2

@@ 4.2 Zero-shot Learning
Recent work such as Xian et al. (2018, 2017); Chang et al. (2008); Larochelle et al. (2008); Palatucci et al. (2009) have shown impressive capability of zero-shot learning in the space of language and vision problems. Recent approaches to zero-shot or few-shot learning for tabular data problems mostly encode tabular data as language, and then leverage large language models (LLMs) for their zero- or few-shot capabilities (see Hegselmann et al. (2023); Nam et al. (2023); Gardner et al. (2024)). These approaches rely on relevant information about the tabular data existing in LLMs – this is most obviously the case when column names are meaningful, but not guaranteed for broad tabular data problems.
@@ 4.3 Adversarial Training
Upon generative adversarial networks (GANs) (Goodfellow et al., 2015; Madry et al., 2018; Kurakin et al., 2017), recent work such as Shafahi et al. (2019) improved on the efficiency by combining the back-propagation steps of the generator and discriminator. However, this method has been shown to suffer from catastrophic overfitting (Andriushchenko & Flammarion, 2020; Kim et al., 2021) without further modifications. Other works focusing on improving the efficiency of GAN training include Wong et al. (2020) and Zhang et al. (2019) where they restrict most of the forward and back propagation within the first layer of the network during adversarial updates. Zhang et al. (2021) in particular noted that weight updates frequently go back and forth in opposite directions in one training epoch, suggesting those updates might be redundant. Many other variations have been introduced to mitigate vanishing gradient and additional challenges of GAN training (Jabbar et al., 2021): failing at finding a Nash-equilibrium (Ratliff et al., 2016), and internal covariate shift (Ioffe, 2015).
@@ 5 Conclusion
In this work, we gave the first effort in exploring the adversarial pre-training of deep zero-shot meta-learners, specifically PFNs. We proposed APT, a zero-shot meta-learner that improves the performance of TabPFN on tabular prediction tasks and matches state-of-the-art GBDTs on small tabular classification tasks. In addition, we proposed a mixture block neural design to eliminate the class size restriction of PFNs, addressing a crucial issue in their generalizability to broad classification problems. As for limitations, APT does not outperform GBDTs on large tabular datasets, and shares the quadratic computational complexity of TabPFN. Hence, extensions of this work could explore means of acquiring data representations in a more inexpensive manner. For example, considerable research in recent years has significantly accelerated the transformer and increased its context length (Wu et al., 2022; Bulatov et al., 2023). It is a worthwhile effort for future research to apply these advancements to APT as well as other PFNs. Besides, future research could extend the mixture block to standard (non-zero-shot) classification settings in light of its ability to generalize and greatly accelerate convergence, which could improve the performance of traditional DNNs on small classification datasets. Mixture block or other alternatives to the dense final layer could also be explored in both standard and zero-shot regression settings, which could have an impact on the inductive bias of DNNs and their underperformance in comparison to GBDTs (Grinsztajn et al., 2022) under certain tabular data nature.
@@ Acknowledgements
We thank Tyler Farnan, Gang Mei, and C. Bayan Bruss for the insightful discussions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Background
In this section, we give a brief introduction to PFNs, and specifically the synthetic data generating mechanism of TabPFN. For a more complete description, see Müller et al. (2021); Hollmann et al. (2022); Nagler (2023). Given training dataset $D^{\mathrm{train}}=(X^{\mathrm{train}},\bm{y}^{\mathrm{train}})$, the goal is to approximate the conditional outcome distribution $y^{\mathrm{test}}\sim p(\cdot|\bm{x}^{\mathrm{test}},D^{\mathrm{train}})$ given a test point $\bm{x}^{\mathrm{test}}$. In the Bayesian framework for supervised learning, the prior of the dataset is a hypothesis of the data generating mechanism $\phi$ drawn from hypothesis space $\Phi$, under which $p(\cdot|\bm{x}^{\mathrm{test}},D^{\mathrm{train}})$ is a posterior predictive distribution (PPD) and can be factorized as follows by the Bayes’ rule:
$p(\cdot|\bm{x}^{\mathrm{test}},D^{\mathrm{train}})$ $=\int_{\phi\in\Phi}p(\cdot|\bm{x}^{\mathrm{test}},\phi)p(\phi|D^{\mathrm{train}})\,d\phi$ (3)
$=\int_{\phi\in\Phi}p(\cdot|\bm{x}^{\mathrm{test}},\phi)\frac{p(\phi)p(D^{\mathrm{train}}|\phi)}{p(D^{\mathrm{train}})}\,d\phi$ (4)
$\propto\int_{\phi\in\Phi}p(\cdot|\bm{x}^{\mathrm{test}},\phi)p(D^{\mathrm{train}}|\phi)p(\phi)\,d\phi.$ (5)
PFNs conduct synthetic prior fitting by defining a family of data generating mechanisms $\Phi$ from which independent samples $\bm{x}_{i}\sim p(\bm{x}_{i})=\mathbb{E}_{p(\phi)}[p(\bm{x}_{i}|\phi)]$ and $y_{i}\sim p(y_{i})=\mathbb{E}_{p(\phi)}[p(y_{i}|\bm{x}_{i},\phi)]$ are drawn to compose feature matrix $(X^{\mathrm{train}},X^{\mathrm{test}})$ and response vector $(\bm{y}^{\mathrm{train}},\bm{y}^{\mathrm{test}})$ of a synthetic dataset $D=D^{\mathrm{train}}\cup D^{\mathrm{test}}$, and use a transformer model $q_{\theta}(\cdot|X^{\mathrm{test}},D^{\mathrm{train}})$ to approximate $p(\cdot|X^{\mathrm{test}},D^{\mathrm{train}})$ by minimizing their expected divergence over the synthetic data distribution
$\mathbb{E}_{p(D^{\mathrm{train}},X^{\mathrm{test}})}\left[\mathrm{KL}\left(p(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},D^{\mathrm{train}}),q_{\theta}(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},D^{\mathrm{train}})\right)\right].$ (6)
Since
$\nabla_{\theta}\mathbb{E}_{p(D^{\mathrm{train}},X^{\mathrm{test}})}\left[\mathrm{KL}\left(p(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},D^{\mathrm{train}}),q_{\theta}(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},D^{\mathrm{train}})\right)\right]$ (7)
$=$ $\nabla_{\theta}\mathbb{E}_{p(D^{\mathrm{train}},X^{\mathrm{test}})}\left[\mathrm{H}\left(p(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},D^{\mathrm{train}}),q_{\theta}(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},D^{\mathrm{train}})\right)\right]$ (8)
$=$ $\nabla_{\theta}\mathbb{E}_{p(D^{\mathrm{train}},D^{\mathrm{test}})}\left[-\log q_{\theta}(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},D^{\mathrm{train}})\right],$ (9)
it is equivalent to minimizing the expected negative log-likelihood loss
$\mathcal{L}(q_{\theta})$ $=\mathbb{E}_{p(D)}\left[-\log q_{\theta}(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},D^{\mathrm{train}})\right].$ (10)
@@ A.2 Limitations
Although there is no theoretical limitation on the number of data PFNs can handle, the transformer architecture does entail significant computation complexity and memory usage for large datasets. Besides, given the nature of dense input embedding layer and dense final prediction layer, there is a theoretical limitation on the number of features and the number of classes that PFNs can handle. The former is less of an issue since feature selections or simply random sampling of features can be performed, and PFNs would still yield ideal performance as shown in McElfresh et al. 2024. The latter is a rather big problem for classification tasks because there is hardly any direct and effective work-around.
Appendix B Handling of Large Datasets
Since expanding TabPFN’s capabilities in handling large datasets is not the focus of this work, we only provide two simple adaptations such that APT can practically handle datasets of this nature. We recommend that practitioners try out the concurrent and future developments in more involved model innovations for large datasets, but in case they do not wish to do so, the following approaches can serve as a baseline.
B.1 Uncertainty-based Batch Aggregation
For datasets with large number of samples, to avoid calculating attention spanning all training data points which results in quadratic order of operations and memory usage with respect to data size, we estimate the PPD with batches drawn from the training set:
$p(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},D^{\mathrm{train}})\approx\int_{b}q_{\theta}(\bm{y}^{\mathrm{test}}|X^{\mathrm{test}},b)\cdot p(b|D^{\mathrm{train}}),$ (11)
which is equivalent to drawing uniform samples from training set $D^{\mathrm{train}}$ and scale the resulting predictions with weights $p(b|D^{\mathrm{train}})$. We cap the batch size at $3000$ in alignment with McElfresh et al. (2024). For classification datasets with number of samples larger than $3000$, we split the training set into batches and weigh the resulting predictions in proportion to the batch size (prediction from the last batch may have less weight than the others). For regression tasks, we parameterize model $q_{\theta}(y^{\mathrm{test}}_{i}|\bm{x}^{\mathrm{test}}_{i},b)$ as Gaussian distribution $(\mu_{\theta}(\bm{x}^{\mathrm{test}}_{i},b),\sigma_{\theta}(\bm{x}^{\mathrm{test}}_{i},b))$ similar to Hollmann et al. (2022), and directly produce the point estimation using the inverse variance estimator in inference time:
$\mathbb{E}[y^{\mathrm{test}}_{i}|\bm{x}^{\mathrm{test}}_{i},D^{\mathrm{train}}]$ $\approx\int_{b}\mathbb{E}_{\theta}[y^{\mathrm{test}}_{i}|\bm{x}^{\mathrm{test}}_{i},b]\cdot p(b|D^{\mathrm{train}})$ (12)
$=\left(\sum_{k}\frac{1}{\sigma_{\theta}^{2}(\bm{x}^{\mathrm{test}}_{i},b_{k})}\right)^{-1}\sum_{k=1}^{N}\frac{1}{\sigma_{\theta}^{2}(\bm{x}^{\mathrm{test}}_{i},b_{k})}\mu_{\theta}(\bm{x}^{\mathrm{test}}_{i},b_{k}).$ (13)
The intuition is that, prediction on each batch is weighted by its uncertainty – more weights are put to the predictions that the model is more certain of, and vice versa.
B.2 Patch-based Feature Embedding
We drew inspiration from Dosovitskiy et al. (2020) and developed a patch-based embedding approach that adapts to datasets with arbitrary number of features. In Hollmann et al. (2022), embeddings of $x$ are acquired by padding or clipping the number of features $d_{k}$ to a certain maximum feature size $d^{*}$, such that $x$ can be fed to a dense feedforward $e_{\theta}:\mathbb{R}^{d^{*}}\rightarrow\mathbb{R}^{d_{\mathrm{model}}}$. Instead, we split features into patches, setting $d^{*}$ as the patch size, and only pad the last patch to $d^{*}$ dimensions if $d\not\equiv 0\pmod{d^{*}}$. After feeding each patch to dense feedforward $e_{\theta}$, we pass them to an attention block with optional relative positional encoding (Su et al., 2021; Press et al., 2021), and average pool across the resulting embeddings of patches. Essentially, this is a half-way approach between using a dense feedforward to embed all features, and using an attention block to tokenize each individual feature. In this way, the embedding block can handle features in a more flexible manner while controlling computational complexity and memory usage.
Appendix C Hyperparameter Settings
@@
Appendix D More Results